What is the Probability of obtaining at least one head when 4 coins are tossed simultaneously?
Answer by Theo(13342)

the probability of getting at least one head is equal to 1 minus the probability of getting zero heads.

the probability of getting 0 heads is equal to 1/2 * 1/2 * 1/2 * 1/2 = 1/16.

the probability of getting at least 1 head is equal to 1 - 1/16 = 15/16.